OpenStudy (anonymous):

the product of two consecutive even integers is 224. Find the integers.

hero (hero):

Hint: If you begin with an even number x, then the next number will be two units from x, hence x + 2 Therefore two consecutive integers will be x + (x+2) If we set this equal to 224, then we'll have x + x + 2 = 224 or simply 2x + 2 = 224 Solve for x

OpenStudy (anonymous):

it's product, so it's actually x^2 +2x =224 224 = 2*2*2*2*2*7 now try how to divide the two group and make two consecutive integer. 16*14

OpenStudy (lgbasallote):

@Hero it seems you have misunderstood the question :) it was looking for the PRODUCT not SUM ^_^ @zhantshen...did you use trial and error?